Use dimensional analysis to check if expression P=1/2^Pv^2 +pgh is correct,given that,p is pressure in pascal,p is density in kgm^-3,v is velocity in m/s,g is acceleration due to gravity in m/s^2 and h is the height in m
1
Expert's answer
2019-03-05T12:28:18-0500

We need to determine the dimensions of each part of equation:


"dim\u2061P=MT^{-2} L^{-1}"

"dim\u2061[1\/2 \u03c1v^2 ] =dim\u2061[1\/2] dim\u2061[\u03c1] dim\u2061[v^2 ]=1(ML^{-3} ) (LT^{-1} )^2=MT^{-2} L^{-1}"


"dim\u2061[\u03c1gh] =dim\u2061[\u03c1]dim\u2061[g]dim\u2061[h]=(ML^{-3} )(LT^{-2})(L)=MT^{-2} L^{-1}"

We have the same result in each case, thus the expression is correct.